# LLM-based Resource-Oriented Intention Inference for Static Resource Detection

### v2

@@ Abstract.
Resource leaks, caused by resources not being released after acquisition, often lead to performance issues and system crashes. Existing static detection techniques rely on mechanical matching of predefined resource acquisition/release APIs and null-checking conditions to find unreleased resources, suffering from both (1) false negatives caused by the incompleteness of predefined resource acquisition/release APIs and (2) false positives caused by the unsoundness of resource reachability validation analysis. To overcome these challenges, we propose InferROI, a novel approach that leverages the exceptional code comprehension capability of large language models (LLMs) to directly infer resource-oriented intentions (acquisition, release, and reachability validation) in code. InferROI first prompts the LLM to infer involved intentions for a given code snippet, and then incorporates a two-stage static analysis approach to check control-flow paths for resource leak detection based on the inferred intentions.
We evaluate the effectiveness of InferROI in both resource-oriented intention inference and resource leak detection. Experimental results demonstrate that InferROI achieves a precision of 74.6% and a recall of 81.8% when inferring resource-oriented intentions for 172 code snippets from the DroidLeaks dataset. Additionally, InferROI covers a significant portion (i.e., 67.9% ) of concerned Android resources listed in the dataset. When applied to 86 bugs from the DroidLeaks dataset, InferROI exhibits both a high bug detection rate (53.5%) and a low false alarm rate (8.1%) compared to eight baseline detectors. Moreover, we apply InferROI to resource leak detection in 100 methods from real-world open-source projects, where InferROI identifies 12 unknown resource leak bugs, with 7 of them being confirmed by developers. Overall, this work demonstrates the promising potential of LLMs in resource leak detection.
conference: ACM SIGSOFT International Symposium on Software Testing and Analysis; 16-20 September, 2024; Vienna, Austriacopyright: none
1. Introduction
@@
Resource leaks, stemming from the failure to release acquired resources (e.g., unclosed file handles), represent critical software defects that can lead to runtime exceptions or program crashes. Such leaks are pervasive in software projects (Ghanavati et al., 2020) and even appear in online code examples, even for some accepted answers in Stack Overflow posts (Zhang et al., 2018).
@@
To address this issue, researchers have proposed various automated techniques (Torlak and Chandra, 2010; Wu et al., 2016b; Kellogg et al., 2021; fin, 2023; inf, 2023) for resource leak detection based on static analysis. These techniques primarily rely on two essential components. The first step involves identifying pairs of resource acquisition APIs and their corresponding resource release APIs; Subsequently, the code is analyzed to verify whether the release API is not appropriately invoked after the acquisition API on the control-flow paths, before the acquired resources become unreachable. For instance, consider the API pair for managing locks: LockManager.acquireLock() represents the resource acquisition API, and LockManager.releaseLock() denotes the corresponding resource release API. A resource leak occurs when releaseLock() is not subsequently called after acquireLock().
@@
However, these existing techniques mainly rely on mechanical matching of predefined resource acquisition/release APIs and null-checking conditions to find unreleased resources, leading to the following two limitations. (1) False negatives from the incompleteness of predefined resource acquisition/release APIs: the effectiveness of current resource leak detection techniques (Torlak and Chandra, 2010; Wu et al., 2016b; Kellogg et al., 2021) hinges on the completeness of predefined API pairs, since they cannot detect resource leaks related to un-predefined resource acquisition/release API pairs. For example, the widely adopted detectors (e.g., FindBugs, Infer, and CodeInspection) are confined to detecting only a small subset of common resource classes in JDK and Android (Liu et al., 2019) by default. Thus, the resource leaks involving emerging or less common resources (e.g., AndroidHttpClient in apache httpclient library) cannot be detected by these tools if the specific acquisition/release APIs for AndroidHttpClient are not additionally configured into these tools. (2) False positives from the unsoundness of resource reachability validation analysis: the accuracy of reachability analysis for resources along control-flow paths significantly influences the occurrence of false alarms in existing detection methods, since unreachable resources would not cause leaks even without being released after acquisition. Approaches like those introduced by Torlak et al. (Torlak and Chandra, 2010) determine resource reachability by scrutinizing null-checking conditions (e.g., cur!=null) within if-statements. However, this approach falls short for certain resources with alternative means of determining reachability, such as through API calls (e.g., !bank.isDisabled()). Consequently, false alarms are triggered for resources that are actually unreachable along specific paths.
@@
In this paper, we propose InferROI, a novel approach that leverages large language models (LLMs) to boost static resource leak detection. Our intuition is that LLMs have been trained on massive code and documentation corpus and thus have exceptional code comprehension capability (OpenAI, 2023; Bubeck et al., 2023; Petroni et al., 2019a; Huang et al., 2022a), which are well-suited for inferring resource-oriented intentions (i.e., the functionalities of resource acquisition/release and resource reachability validation) in any given code snippet. InferROI consists of two stages, i.e., resource-oriented intention inference and lightweight resource leak detection. (1) For a given code snippet, InferROI first prompts the LLM to directly infer the involved resource acquisition, release, and reachability validation intentions, without relying on any prior knowledge of resource API pairs. The intentions generated by LLMs (which are described in natural language) are then formalized into formal expressions, so as to facilitate the subsequent static analysis stage. (2) InferROI then incorporates a two-stage static analysis approach to check control-flow paths for resource leaks based on the inferred intentions. In particular, InferROI first identifies potential leak-risky paths based on the inferred resource acquisition/release APIs and then prunes leak-risky paths with unreachable resources based on the inferred reachability validation, so as to reduce false alarms of resource leak detection.
@@
While our approach is general and not limited to specific programming languages, we currently implement InferROI for Java given its prevalence. We perform an extensive evaluation to investigate the effectiveness of InferROI in inferring resource-oriented resources and detecting resource leaks. First, we collect 86 bugs in the DroidLeaks dataset (Liu et al., 2019), producing 172 (2 $\times$ 86) code snippets (buggy version and fixed version). We manually annotate the resource-oriented intentions involved in the code snippets, and check how many are correctly inferred by InferROI. The experimental results show that InferROI achieves a precision of 74.6% and a recall of 81.8% in intention inference, covering 19 (i.e., 67.9% ) concerned Android resources listed in DroidLeaks. Second, we apply InferROI on the 172 code snippets to evaluate its effectiveness in resource leak detection. On the 86 buggy-version code snippets, InferROI achieves a relatively high bug detection rate (53.5%), compared to 3.4%~56.9% by the baselines; On the 86 fixed-version code snippets, InferROI generates a low false alarm rate (8.1%), comparing to 0~41.8% by the baselines. Third, we apply InferROI on resource detection for real-world open-source projects. The experimental results show that in 100 methods, InferROI detects 12 unknown resource leaks, among which 7 have been confirmed by developers as of the submission time. Finally, we conduct an ablation study to investigate the contributions of the two key components and the results demonstrate the importance of combining the LLM-based inference and the static analysis. Overall, our work shows the promising potential of LLMs in resource leak detection.
@@ 2. Motivating Example
Figure 1 illustrates the code diff of a commit (cod, 2023a) of fixing a resource leak associated with AndroidHttpClient, as observed in the DroidLeaks dataset. In the buggy version, the API AndroidHttpClient.newInstance(...) is called to acquire an HTTP client resource at line 166, but the corresponding release API AndroidHttpClient.close() is not called, leading to a resource leak. In the fixed version, a finally structure is introduced at lines 184-187 to ensure the proper release of the acquired HTTP client resource. Through this example, we demonstrate the following key insights:
Figure 1. Comparison between Buggy Version (Left) and Fixed Version (Right) Associated with AndroidHttpClient Leak
1. Identifying resource acquisition/release more comprehensively. In the aforementioned example, the presence of a resource leak related to the AndroidHttpClient class goes unnoticed by all the eight subject detectors (i.e., Code Inspection (cod, 2023b), Infer (inf, 2023), Lint (lin, 2023), FindBugs (fin, 2023), Relda2-FS (Wu et al., 2016a), RElda2-FI (Wu et al., 2016a), Elite (Liu et al., 2016), and Verifier (Vekris et al., 2012)) evaluated in the DroidLeaks dataset. The reason for this oversight lies in the fact that AndroidHttpClient is not included in the predefined API sets used by these detectors. This limitation highlights a fundamental constraint of the existing detection paradigm, which primarily relies on locating API calls for resource acquisition/release through the resolution of API signatures in code (i.e., type inference) and subsequent matching with predefined API names.
@@
Furthermore, in addition to AndroidHttpClient, a multitude of APIs are available in various libraries/packages that implement the functionality of HTTP client. For instance, within the selenium library (sel, 2023), there exist related APIs like HttpClient, JdkHttpClient, NettyClient, TracedHttpClient, NettyDomainSocketClient, PassthroughHttpClient, among others. The eight detectors evaluated in DroidLeaks also fail to detect resource leaks across all of them. This failure can be attributed to the detectors’ limited capability of capturing the underlying common characteristics shared by these APIs, which essentially represent the same resource concept, namely, an HTTP client, and its usage. The presence of this limitation further emphasizes the necessity of incorporating background knowledge and code understanding in resource acquisition and release identification.
@@
Existing detection techniques, such as Torlak et al. (Torlak and Chandra, 2010) and Cherem et al. (Cherem et al., 2007), typically identify reachability validation by matching null-checking conditions. While this approach provides a simple and intuitive means to recognize the reachability validation at line 185, it may overlook certain validation intentions present in other code. For specific resource types, dedicated APIs, such as isClosed() and isDisabled(), are employed to perform reachability validation. In a similar manner to resource acquisition/release identification, a more comprehensive identification of such validation intentions necessitates the incorporation of background knowledge pertaining to common validation verbs, in conjunction with a deep understanding of the code contexts, such as the if-conditions. By considering these elements, resource reachability validation can be assessed more comprehensively, enabling a more accurate detection of resource leaks in code.
3. Approach
Based on these insights, we propose InferROI, a novel approach that directly Infers Resource-Oriented Intentions in code to boost static resource leak detection, by leveraging the power of large language models (LLMs). In the approach, LLMs function as extensive knowledge bases with exceptional code understanding capabilities (OpenAI, 2023; Bubeck et al., 2023; Petroni et al., 2019a; Huang et al., 2022a), rendering them well-suited for inferring resource-oriented intentions.
@@ 3. Approach
Figure 2. Approach Overview of InferROI
3.1. Resource-Oriented Intentions
@@
Resource Acquisition. Statements acquire resources (e.g., locks or database connections), which are normally API invocations and involved with objects/variables of resources (e.g., lock or db). To formally express the intentions of resource acquisition, we introduce the notation $\bm{ACQUIRE(var,lineno)}$, where $var$ denotes the variable storing the acquired resource, and $lineno$ indicates the line number containing the statement responsible for acquiring the resource.
Resource Release. Likewise, statements release previously acquired resources. We formalize the intentions of resource release as $\bm{RELEASE(var,lineno)}$, where $var$ corresponds to the variable storing the resource to be released, and $lineno$ indicates the line number of the statement responsible for performing the release operation.
Resource Reachability Validation. Conditional statements aim to validate whether acquired resources remain reachable within specific branches. To formalize the intentions of resource reachability validation, we employ the notation $\bm{VALIDATE(var,lineno)}$, where $var$ represents the variable storing the resource being validated, and $lineno$ indicates the line number of the statement responsible for performing the validation operation.
Example 3.1. In the method depicted in Figure 1, the intentions of the API calls at line 167, which aims to acquire a HTTP client resource, can be represented as $ACQUIRE(\texttt{\small client},167)$. The intention of the API call at line 186, responsible for releasing the acquired HTTP client resource, can be denoted as $RELEASE(\texttt{\small client},186)$. The intention of the if statement at line 185, intended to validate the reachability of the acquired HTTP client resource, can be expressed as $VALIDATE(\texttt{\small client},185)$.
3.2. LLM-Based Intention Inference
@@
3.2.1. Feeding LLM with Prompt
For a given code, we construct a prompt using the template depicted in Figure 3. This prompt is then fed into the LLM (i.e., GPT-4 used in this work) for intention inference. The prompt template is designed following the best practices for prompt engineering by OpenAI (ope, 2023), consisting of three main components: Task Description & Instructions, Output Format Specification, and Code Placeholder.
Figure 3. Prompt Template
@@ 3.2.1. Feeding LLM with Prompt
Example 3.2. Applying the template, we construct a prompt for the fixed version (right) illustrated in Figure 1. We then employ the gpt-4 model (gpt, 2023) to generate an answer, as shown in Figure 4. Notably, the gpt-4 model exhibits the capability to effectively analyze code semantics and discern resource-oriented intentions. Within the generated answer, resource acquisition intentions are highlighted in red , resource release intentions in green , and resource reachability validation intentions in cyan .
Figure 4. Answer Generated by gpt-4 Model
3.2.2. Extracting Formal Expressions
@@
Example 3.3. From the answer illustrated in Figure 4, the three highlighted lines are matched by the matching patterns. Subsequently, the three formal intention expressions are extracted and instantiated as $ACQUIRE(\texttt{\small client},167)$, $RELEASE(\texttt{\small client},186)$, and $VALIDATE(\texttt{\small client},185)$.
3.3. Static Resource Leak Detection

### v3

# Inferring Resource-Oriented Intentions using LLMs for Static Resource Leak Detection

@@ Abstract
Resource leaks, caused by resources not being released after acquisition, often lead to performance issues and system crashes. Existing static detection techniques rely on mechanical matching of predefined resource acquisition/release APIs and null-checking conditions to find unreleased resources, suffering from both (1) false negatives caused by the incompleteness of predefined resource acquisition/release APIs and (2) false positives caused by the incompleteness of resource reachability validation identification.
To overcome these challenges, we propose InferROI, a novel approach that leverages the exceptional code comprehension capability of large language models (LLMs) to directly infer resource-oriented intentions (acquisition, release, and reachability validation) in code. InferROI first prompts the LLM to infer involved intentions for a given code snippet, and then incorporates a two-stage static analysis approach to check control-flow paths for resource leak detection based on the inferred intentions. We evaluate the effectiveness of InferROI in both resource-oriented intention inference and resource leak detection. Experimental results on the DroidLeaks and JLeaks datasets demonstrate InferROI achieves promising bug detection rate (59.3% and 64.8%) and false alarm rate (18.6% and 24.0%). Compared to three industrial static detectors, InferROI detects 14~45 and 167~503 more bugs in DroidLeaks and JLeaks, respectively. When applied to real-world open-source projects, InferROI identifies 26 unknown resource leak bugs, with 7 of them being confirmed by developers. Finally, manual annotation indicated that InferROI achieved a precision of 74.6% and a recall of 81.8% in intention inference, covering more than 60% resource types involved in the datasets. The results of an ablation study underscores the importance of combining LLM-based inference with static analysis.
I Introduction
@@
Resource leaks, stemming from the failure to release acquired resources (e.g., unclosed file handles), represent critical software defects that can lead to runtime exceptions or program crashes. Such leaks are pervasive in software projects [1] and even appear in online code examples, even for some accepted answers in Stack Overflow posts [2].
@@
To address this issue, researchers have proposed various automated techniques [3, 4, 5, 6, 7] for resource leak detection based on static analysis. These techniques primarily rely on two essential components. The first step involves identifying pairs of resource acquisition APIs and their corresponding resource release APIs; Subsequently, the code is analyzed to verify whether the release API is not appropriately invoked after the acquisition API on the control-flow paths, before the acquired resources become unreachable. For instance, consider the API pair for managing locks: LockManager.acquireLock() represents the resource acquisition API, and LockManager.releaseLock() denotes the corresponding resource release API. A resource leak occurs when releaseLock() is not subsequently called after acquireLock().
@@
However, these existing techniques mainly rely on mechanical matching of predefined resource acquisition/release APIs and null-checking conditions to find unreleased resources, leading to the following two limitations. (1) False negatives from the incompleteness of predefined resource acquisition/release APIs: the effectiveness of current resource leak detection techniques [3, 4, 5] hinges on the completeness of predefined API pairs, since they cannot detect resource leaks related to un-predefined resource acquisition/release API pairs. For example, the widely adopted detectors (e.g., SpotBugs, Infer, and CodeInspection) are confined to detecting only a small subset of common resource classes in JDK and Android [8] by default. Thus, the resource leaks involving emerging or less common resources (e.g., AndroidHttpClient in apache httpclient library) cannot be detected by these tools if the specific acquisition/release APIs for AndroidHttpClient are not additionally configured into these tools. (2) False positives from the incompleteness of resource reachability validation identification: the accuracy of reachability analysis for resources along control-flow paths significantly influences the occurrence of false alarms in existing detection methods, since unreachable resources would not cause leaks even without being released after acquisition. Approaches like those introduced by Torlak et al. [3] determine resource reachability by scrutinizing null-checking conditions (e.g., cur!=null) within if-statements. However, this approach falls short for certain resources with alternative means of determining reachability, such as through API calls (e.g., !bank.isDisabled()). Consequently, false alarms are triggered for resources that are actually unreachable along specific paths.
@@
In this paper, we propose InferROI, a novel approach that leverages large language models (LLMs) to boost static resource leak detection. Our intuition is that LLMs have been trained on massive code and documentation corpus and thus have exceptional code comprehension capability [9, 10, 11, 12], suggesting their potential for inferring resource-oriented intentions (i.e., the functionalities of resource acquisition/release and resource reachability validation) in any given code snippet. InferROI consists of two stages, i.e., resource-oriented intention inference and lightweight resource leak detection. (1) For a given code snippet, InferROI first prompts the LLM to directly infer the involved resource acquisition, release, and reachability validation intentions, without relying on any prior knowledge of resource API pairs. The intentions generated by LLMs (which are described in natural language) are then formalized into formal expressions, so as to facilitate the subsequent static analysis stage. (2) InferROI then incorporates a two-stage static analysis approach to check control-flow paths for resource leaks based on the inferred intentions. In particular, InferROI first identifies potential leak-risky paths based on the inferred resource acquisition/release APIs and then prunes leak-risky paths with unreachable resources based on the inferred reachability validation, so as to reduce false alarms of resource leak detection.
@@
While our approach is general and not limited to specific programming languages, we currently implement InferROI for Java given its prevalence. We conduct an extensive evaluation to assess the efficacy of InferROI in detecting resource leaks. Firstly, we apply InferROI to 86 and 784 bugs sourced from the DroidLeaks [8] and JLeaks [13] datasets, respectively, to evaluate its effectiveness in resource leak detection. The results reveal InferROI demonstrates a bug detection rate of 59.3% with a false alarm rate of 18.6% for DroidLeaks, and a bug detection rate of 64.8% with a false alarm rate of 24.0% for JLeaks. Compared to three industrial static detectors, InferROI detects 14~45 and 167~503 more bugs in DroidLeaks and JLeaks, respectively, while maintaining a comparable false alarm rate. Secondly, we employ InferROI for resource detection in real-world open-source projects. Our experiments on 100 methods uncover 26 previously unknown resource leaks, 7 of which are confirmed by developers as of the submission time. Finally, manual annotation for resource-oriented intentions indicated that InferROI achieved a precision of 74.6% and a recall of 81.8% in intention inference, covering 19 (67.9%) and 221 (60.1%) resource types involved in DroidLeaks and JLeaks, respectively. The results of an ablation study underscores the importance of combining LLM-based inference with static analysis.
@@ II Motivating Example
Figure 1: Comparison between Buggy Version (Left) and Fixed Version (Right) Associated with AndroidHttpClient Leak
Figure 1 illustrates the code diff of a commit [14] of fixing a resource leak associated with AndroidHttpClient, as observed in the DroidLeaks dataset. In the buggy version, the API AndroidHttpClient.newInstance(...) is called to acquire an HTTP client resource at line 166, but the corresponding release API AndroidHttpClient.close() is not called, leading to a resource leak. In the fixed version, a finally structure is introduced at lines 184-187 to ensure the proper release of the acquired HTTP client resource. Through this example, we demonstrate the following key insights:
1. Identifying resource acquisition/release more comprehensively. In the aforementioned example, the presence of a resource leak related to the AndroidHttpClient class goes unnoticed by all the eight subject detectors (i.e., Code Inspection [15], Infer [7], Lint [16], SpotBugs [6], Relda2-FS [17], RElda2-FI [17], Elite [18], and Verifier [19]) evaluated in the DroidLeaks dataset. The reason for this oversight lies in the fact that AndroidHttpClient is not included in the predefined API sets used by these detectors. This limitation highlights a fundamental constraint of the existing detection paradigm, which primarily relies on locating API calls for resource acquisition/release through the resolution of API signatures in code (i.e., type inference) and subsequent matching with predefined API names.
@@
Furthermore, in addition to AndroidHttpClient, a multitude of APIs are available in various libraries/packages that implement the functionality of HTTP client. For instance, within the selenium library [20], there exist related APIs like HttpClient, JdkHttpClient, NettyClient, TracedHttpClient, NettyDomainSocketClient, PassthroughHttpClient, among others. The eight detectors evaluated in DroidLeaks also fail to detect resource leaks across all of them. This failure can be attributed to the detectors’ limited capability of capturing the underlying common characteristics shared by these APIs, which essentially represent the same resource concept, namely, an HTTP client, and its usage. The presence of this limitation further emphasizes the necessity of incorporating background knowledge and code understanding in resource acquisition and release identification.
@@
Existing detection techniques, such as Torlak et al. [3] and Cherem et al. [21], typically identify reachability validation by matching null-checking conditions. While this approach provides a simple and intuitive means to recognize the reachability validation at line 185, it may overlook certain validation intentions present in other code. For specific resource types, dedicated APIs, such as isClosed() and isDisabled(), are employed to perform reachability validation. In a similar manner to resource acquisition/release identification, a more comprehensive identification of such validation intentions necessitates the incorporation of background knowledge pertaining to common validation verbs, in conjunction with a deep understanding of the code contexts, such as the if-conditions. By considering these elements, resource reachability validation can be assessed more comprehensively, enabling a more accurate detection of resource leaks in code.
III Approach
Based on these insights, we propose InferROI, a novel approach that directly Infers Resource-Oriented Intentions in code to boost static resource leak detection, by leveraging the power of large language models (LLMs). In the approach, LLMs function as extensive knowledge bases with exceptional code understanding capabilities [9, 10, 11, 12], rendering them well-suited for inferring resource-oriented intentions.
@@ III Approach
Figure 2: Approach Overview of InferROI
III-A Resource-Oriented Intentions
@@
Resource Acquisition. Statements acquire resources (e.g., locks or database connections), which are normally API invocations and involved with objects/variables of resources (e.g., lock or db). To formally express the intentions of resource acquisition, we introduce the notation $\bm{ACQ(var,lineno)}$, where $var$ denotes the variable storing the acquired resource, and $lineno$ indicates the line number containing the statement responsible for acquiring the resource.
Resource Release. Likewise, statements release previously acquired resources. We formalize the intentions of resource release as $\bm{REL(var,lineno)}$, where $var$ corresponds to the variable storing the resource to be released, and $lineno$ indicates the line number of the statement responsible for performing the release operation.
Resource Reachability Validation. Conditional statements aim to validate whether acquired resources remain reachable within specific branches. To formalize the intentions of resource reachability validation, we employ the notation $\bm{VAL(var,lineno)}$, where $var$ represents the variable storing the resource being validated, and $lineno$ indicates the line number of the statement responsible for performing the validation operation.
Example 3.1. In the method depicted in Figure 1, the intentions of the API calls at line 167, which aims to acquire a HTTP client resource, can be represented as $ACQ({\texttt{\footnotesize client}},167)$. The intention of the API call at line 186, responsible for releasing the acquired HTTP client resource, can be denoted as $REL({\texttt{\footnotesize client}},186)$. The intention of the if statement at line 185, intended to validate the reachability of the acquired HTTP client resource, can be expressed as $VAL({\texttt{\footnotesize client}},185)$.
III-B LLM-Based Intention Inference
@@
III-B1 Feeding LLM with Prompt
For a given code, we construct a prompt using the template depicted in Figure 3. This prompt is then fed into the LLM (i.e., GPT-4 used in this work) for intention inference. The prompt template is designed following the best practices for prompt engineering by OpenAI [22], consisting of three main components: Task Description & Instructions, Output Format Specification, and Code Placeholder.
Figure 3: Prompt Template
@@ III-B1 Feeding LLM with Prompt
Example 3.2. Applying the template, we construct a prompt for the fixed version (right) illustrated in Figure 1. We then employ the GPT-4 model [23] to generate an answer, as shown in Figure 4. Notably, the GPT-4 model exhibits the capability to effectively analyze code semantics and discern resource-oriented intentions. Within the generated answer, resource acquisition intentions are highlighted in red , resource release intentions in green , and resource reachability validation intentions in cyan .
Figure 4: Answer Generated by GPT-4 Model
III-B2 Extracting Formal Expressions
@@
Example 3.3. From the answer illustrated in Figure 4, the three highlighted lines are matched by the matching patterns. Subsequently, the three formal intention expressions are extracted and instantiated as $ACQ({\texttt{\footnotesize client}},167)$, $REL({\texttt{\footnotesize client}},186)$, and $VAL({\texttt{\footnotesize client}},185)$.
III-C Static Resource Leak Detection